OpenStudy (anonymous):

Solve: sinx+cosx+sin2x>1

OpenStudy (anonymous):

i imagine it gon be a hell of a lengthy work :D

OpenStudy (anonymous):

I know that sin2x=2sinxcosx, but i cannot facotorize this on right way :(

OpenStudy (anonymous):

just graph it. Save tones of times :DD

OpenStudy (anonymous):

Yeah and wolfram saves time, but i need procedure :)

OpenStudy (dumbcow):

yeah this wont factor at all you just have to recognize that at x=0 you get 1 and at x=pi/2 you get 1 so its greater than 1 in the interval (0,pi/2)

OpenStudy (anonymous):

and how i know that function f:R->R f(x)=sinx+cosx+sin2x doesnt intersect x ray between this segment?

OpenStudy (ranga):

sin(x) + cos(x) + sin(2x) >1 sin(x) + cos(x) + 2sin(x)cos(x) >1 Replace the 1 on the right with sin^2(x) + cos^2(x) sin(x) + cos(x) + 2sin(x)cos(x) > sin^2(x) + cos^2(x) sin(x) + cos(x) > sin^2(x) + cos^2(x) - 2sin(x)cos(x) sin(x) + cos(x) > {(sin(x) - cos(x)}^2 The RHS being a square is always >= 0 So we need to solve for sin(x) + cos(x) > 0

OpenStudy (ranga):

Maybe the conclusion in the last line is not correct.

OpenStudy (anonymous):

@ranga I was gonna say so :DD

OpenStudy (ranga):

But if I plot sin(x) + cos(x) + sin(2x); y = 1 and sin(x) + cos(x) the three graphs nicely intersect at all the points where the original function is > 1.

OpenStudy (anonymous):

:ranga it isnt correct beacause for 2pi/3 sin(x) + cos(x) > 0, but sin(x) + cos(x) + sin(2x) >1 is not correct...

OpenStudy (anonymous):

well, consider x > x^2. The solution isn't x > 0 simply because x^2 > 0

OpenStudy (ranga):

0 <= ( sin(x) - cos(x) )^2 <= 2 can this fact be used?

OpenStudy (anonymous):

idk :D why is that needed? :D

OpenStudy (ranga):

because sin(x) + cos(x) > {(sin(x) - cos(x)}^2

OpenStudy (mertsj):

Is that sin^2x or sin(2x)?

OpenStudy (anonymous):

sin(2x)

OpenStudy (ranga):

Wherever sin(x) + cos(x) > 1 so is the original function.

OpenStudy (anonymous):

how i know that? :D

OpenStudy (anonymous):

without looking graph...

OpenStudy (ranga):

I am reading from the graph and trying to see if that fact can be used as a hint.

OpenStudy (anonymous):

your solution indeed matches with the solution provided by wolfram alpha but I'm still trying to figure out how 0 <= ( sin(x) - cos(x) )^2 <= 2 is good fact to use

OpenStudy (anonymous):

i got it... substitution x=pi/4-y solves it :)